# On the nature of buckling instability in galactic bars

### v1

@@ 5 Making pretzels from bananas
Let us consider a few examples of horizontal frequencies of stellar orbits after buckling, $f_{x}^{\prime}$, from the range shown in the right panel of Fig. 11. The selected values are listed in the first column of Table 1. Using equation (1) one can immediately find the corresponding vertical frequencies after buckling, $f_{z}^{\prime}$. These values are given in the second column of the Table and their ratios, $f_{z}^{\prime}/f_{x}^{\prime}$, in the third. The range of values of these ratios, between 1.5 and 2, covers the whole range seen in the red histogram in Fig. 2. Given that during buckling $f_{z}^{\prime}=2f_{x}$ one can immediately recover the initial horizontal frequencies, $f_{x}$ (fourth column). The typical amplitudes of the oscillations along the bar after and before buckling, giving an estimate of the size of the stellar orbits in this dimension, $a_{x}^{\prime}$ and $a_{x}$, can then be found from fitting the very tight relations $f_{x}^{\prime}(a_{x}^{\prime})$ and $f_{x}(a_{x})$ in Fig. 11. Their values are given in the fifth and sixth column of Table 1.
Figure 13: Examples of pretzel orbits formed as a result of buckling, viewed edge-on. The banana orbit is included in the lower panel. The frequency ratios of the orbits are $f_{z}^{\prime}/f_{x}^{\prime}=3/2$, 5/3, 7/4, 9/5, and 2, from top to bottom.
The passage of the bending wave may be interpreted as speeding up the motion of a star in the horizontal plane. For example (all frequencies in units of Gyr-1), a star with $f_{x}=7$ made 7 cycles in 1 Gyr and after buckling it makes 8 due to the bending wave: $f_{x}^{\prime}=f_{x}+(1/7)f_{x}$, a star with $f_{x}=25/3=8.33$ now has $f_{x}^{\prime}=25/3+(1/5)25/3=10$, while a star with $f_{x}=15$ makes 5 more cycles and has $f_{x}^{\prime}=f_{x}+(1/3)f_{x}=20$. These regularities can be generalized to
@@
where $f_{\rm p}$, as always, is the pattern speed of the bar. The formula makes clear the special significance of $f_{x}=f_{x}^{\prime}=2f_{\rm p}$, because $\psi=0$ for this particular frequency. It means that the stellar orbits with frequencies equal to twice the pattern speed of the bar do not get modified after buckling (last row of Table 1).
@@
where $a,b,c$ are the amplitudes of the oscillations along $x,y,z$, respectively, and $m$ is the ratio of vertical to horizontal frequency. It is assumed here that in the horizontal plane $xy$ the orbits were and remain ellipses, that is they originate from the x1 orbits well-known from the classical studies of the orbital structure of bars. In this case, $f_{y}=f_{x}$ before buckling and $f_{y}^{\prime}=f_{x}^{\prime}$ after buckling. On the other hand, the ratio between the vertical and horizontal frequency varies during buckling and changes from $f_{z}/f_{x}>2$ before the event to $f_{z}^{\prime}/f_{x}^{\prime}\leq 2$ afterwards, as discussed above. For resonant orbits, $f_{z}^{\prime}=2f_{x}$ so a banana orbit is obtained for $m=2$ in equation (6). With the constant $\phi_{0}=0$, varying $\phi$ from 0 to $2\pi$ we obtain the “pure” banana shown in the lower panel of Fig. 13. For this plot I assumed $a=a_{x}^{\prime}=7.3$ kpc (in agreement with the values from Table 1 for $f_{x}=f_{x}^{\prime}=5$ Gyr-1) and $b=c=a/2$. Adopting a non-zero value of $\phi_{0}$ shifts the banana orbit toward an $\infty$-like orbit, for example the orbit shown in Fig. 6 was obtained with $\phi_{0}=-\pi/5$.
At the end of the first phase of buckling all orbits can be approximated as banana-like orbits with $f_{z}^{\prime}=2f_{x}$. Using equation (4) one gets a relation between the frequencies after buckling
@@
where $n=(1+\psi)/(2\psi)$ and the values of the parameter $n$ for the selected orbits are given in the last column of Table 1. In order to obtain the new orbits from the Lissajous curves given in the parametric form (6) one thus needs to replace $m=2$ valid for the banana orbits with $(2-1/n)$. These values are obviously equal to the $f_{z}^{\prime}/f_{x}^{\prime}$ ratios given in the third column of Table 1. The orbits obtained in this way using the Lissajous curves are shown in Fig. 13. The sizes of the orbits along $x$ were adjusted to the values of $a=a_{x}^{\prime}$ from Table 1 while $b=c=a/2$ and $\phi_{0}=0$ were kept for all of them. It should be noted that in order to close, the orbits require to cover $4\pi$, $6\pi$, $8\pi$, and $10\pi$ in $\phi$ for $n=2,3,4$ and 5, respectively.
The origin of these pretzel-like orbits can be understood recalling that the bending wave speeds up the motion in the horizontal plane. Let us focus on the first example of a fish-like orbit shown in the upper panel of Fig. 13. Starting at the upper right corner of the box, at the highest $z$ position, and following the orbit one can see that indeed, in comparison with the original banana orbit (bottom panel) the star reaches the next maximum of $z$ not after half of a cycle, but much further along the horizontal angle. Overall, the result is that the star makes only 3 full oscillations along $z$ for each 2 oscillations in $x$, as expected from the frequency ratio for this orbit, $f_{z}^{\prime}/f_{x}^{\prime}=3/2$. A similar behavior is observed for the next orbits shown in Fig. 13, only to a lesser extent. The orbits with $f_{z}^{\prime}/f_{x}^{\prime}=$ 5/3, 7/4, 9/5 make 5, 7 and 9 oscillations in $z$ for each 3, 4 and 5 oscillations in $x$, respectively. The limiting case of this sequence is obviously the banana orbit with $f_{z}^{\prime}/f_{x}^{\prime}=2$ which remains unchanged.
The positions of the first four orbits from Table 1 and Fig. 13 in the $a_{x}^{\prime}(a_{x})$ and $f_{x}^{\prime}(f_{x})$ planes are shown as black dots in Fig. 12. The variation of these two quantities in the second phase of buckling suggests that they change adiabatically. As discussed in section 3.6.2 of Binney & Tremaine (2008), the increase of the horizontal frequency of a harmonic oscillator by a factor $(1+\psi)$ described by equation (4) should correspond to the decrease of the amplitude of oscillations according to $a_{x}^{\prime}=a_{x}(1+\psi)^{-1/2}$.
@@ 6 Discussion
I this work I revisited the issue of the origin and interpretation of the buckling instability in galactic bars. I proposed that the phenomenon can be divided into two distinct phases. The first one involves the growth of the distortion of the bar out of the disk plane due to the vertical resonance of stellar orbits in the bar. The growth of the distortion can be described with the mechanism of the driven harmonic oscillator where the initial small perturbation acts as a driving force and modifies the vertical frequencies of other orbits. The distortion retains its orientation with respect to the bar until reaching a maximum value.
@@
As a result of the increased horizontal frequencies, in the second phase of buckling the pattern speed of the vertical distortion is no longer equal to the pattern speed of the bar, but higher, and the distortion starts to wind up. The banana orbits that dominated the orbital structure of the bar at the time of the resonance evolve into pretzel-like orbits with vertical-to-horizontal frequency ratios below 2. The winding up of the distortion also explains why the bar becomes weaker after buckling. Namely, the winding up is associated with faster rotation of the stars in the horizontal plane which means that the orbits become less radial.
@@
The equation of driven harmonic oscillator was previously used in the context of bar buckling by Collier (2020) in an attempt to relate buckling to the phenomenon of Landau damping. Here instead, this equation was used to demonstrate how the initial small deviation from the vertical symmetry can build up leading to a strong global distortion of the bar and the scenario bears no relation to Landau damping.
@@
The violent buckling event seen in the simulation described here, leading to the strong changes in the orbital structure of the bar, seems to be characteristic of pure disks embedded in dark matter halos, with no bulge component. In a recent study, McClure et al. (2024) looked into the phenomenon using simulations with different bulges of increasing mass. In the case without the bulge, their results are very similar to the ones discussed here, namely, after buckling the orbital frequencies display a very tight relation similar to the one shown in the right panel of Fig. 1. With increasing bulge mass, the resonant orbits still appear but they do not evolve strongly toward vertical-to-horizontal frequency ratios below 2, but remain close to this resonant value. It seems therefore that the bulge does not inhibit the appearance of the vertical resonance but it does prevent the passage to the second phase of buckling as described here. This may be understandable given the importance of mass redistribution in changing the horizontal frequencies, mentioned above. With the additional massive component in the center this must certainly look different.
@@
Recently, Zozulia et al. (2024) studied the vertical evolution of a mature simulated bar, long after the violent buckling event. They reported behavior similar to the one described here, namely the catching of the orbits at resonance and their later passage to vertical-to-horizontal frequency ratios below 2, which they refer to as the “circulation mode”. This stage of bar evolution corresponds to the period after $t=5$ Gyr in the simulation discussed here when the distortion is preserved only in the outer part of the bar (last panels of Fig. 4). As discussed in L19, this distortion survives for a few Gyr, shifting toward the outer part of the bar and seeding the second episode of buckling later on. It therefore seems that as the bar grows, its outer parts continuously increase amplitudes of their oscillations in the vertical direction in a kind of quiescent buckling. The process described by Zozulia et al. (2024) is thus not essentially different from the violent buckling event studied here. The latter only happens on a shorter timescale and in the whole inner part of the bar.
@@
A few open questions remain concerning the analysis presented here. For example, it is still not clear why the pattern speed of the vertical distortion increases in such a particular way in the second stage of buckling and becomes equal to a third of the circular frequency. A related question is why the banana orbits in the outer part of the bar remain unchanged. This region is characterized by the horizontal frequency equal to twice the pattern speed of the bar so that the pattern speed of the vertical distortion remains equal to the pattern speed of the bar and the distortion does not wind up there.
@@
In summary, this work demonstrates that the phenomenon of buckling instability is not related to the fire-hose instability known from plasma physics but instead is triggered by the vertical resonance of bar orbits. The resonance leads to the growing distortion of the bar which later winds up resulting in a transformation of banana-like orbits into pretzel-like ones and the formation of a BP shape.
@@ Acknowledgements.
Computations for this work have been performed using the computer cluster at the Nicolaus Copernicus Astronomical Center of the Polish Academy of Sciences (CAMK PAN).

### v2

@@ 5 Making pretzels from bananas
I now consider a few examples of horizontal frequencies of stellar orbits after buckling, $f_{x}^{\prime}$, from the range shown in the right panel of Fig. 10. The selected values are listed in the first column of Table 1. Using Equation (1), one can immediately find the corresponding vertical frequencies after buckling, $f_{z}^{\prime}$. These values are given in the second column of the table, and their ratios, $f_{z}^{\prime}/f_{x}^{\prime}$, are given in the third. The range of values of these ratios, which is between 1.5 and 2, covers the whole set seen in the red histogram in Fig. 2. Given that during buckling $f_{z}^{\prime}=2f_{x},$ one can immediately recover the initial horizontal frequencies, $f_{x}$ (fourth column). The typical amplitudes of the oscillations along the bar after and before buckling, giving an estimate of the size of the stellar orbits in this dimension, $a_{x}^{\prime}$ and $a_{x}$, can then be found from fitting the very tight relations $f_{x}^{\prime}(a_{x}^{\prime})$ and $f_{x}(a_{x})$ in Fig. 10. Their values are given in the fifth and sixth columns of Table 1.
Figure 12: Examples of pretzel orbits formed as result of buckling, viewed edge-on. The banana orbit is included in the lower panel. The frequency ratios of the orbits are $f_{z}^{\prime}/f_{x}^{\prime}=3/2$, 5/3, 7/4, 9/5, and 2, from top to bottom.
The passage of the bending wave may be interpreted as speeding up the motion of a star in the horizontal plane. For example (all frequencies in units of 1 Gyr-1), a star with $f_{x}=7$ made seven cycles in 1 Gyr, and after buckling it makes eight due to the bending wave: $f_{x}^{\prime}=f_{x}+(1/7)f_{x}$; a star with $f_{x}=25/3=8.33$ now has $f_{x}^{\prime}=25/3+(1/5)25/3=10$, while a star with $f_{x}=15$ makes five more cycles and has $f_{x}^{\prime}=f_{x}+(1/3)f_{x}=20$. These regularities can be generalized to
@@
where $f_{\rm p}$, as always, is the pattern speed of the bar. The formula makes the special significance of $f_{x}=f_{x}^{\prime}=2f_{\rm p}$ clear, because $\psi=0$ for this particular frequency. It means that the stellar orbits with frequencies equal to twice the pattern speed of the bar are not modified after buckling (last row of Table 1).
@@
where $a,b,c$ are the amplitudes of the oscillations along $x,y,z$, respectively, and $m$ is the ratio of the vertical and the horizontal frequency. It is assumed here that in the horizontal plane $xy$ the orbits were, and remain, ellipses; that is, they originate from the x1 orbits. In this case, $f_{y}=f_{x}$ before buckling and $f_{y}^{\prime}=f_{x}^{\prime}$ after buckling. On the other hand, the ratio of the vertical and the horizontal frequency varies during buckling and changes from $f_{z}/f_{x}>2$ before the event to $f_{z}^{\prime}/f_{x}^{\prime}\leq 2$ afterwards, as discussed above. For resonant orbits, $f_{z}^{\prime}=2f_{x},$ so a banana orbit is obtained for $m=2$ in Equation (5). With the constant $\phi_{0}=0$, varying $\phi$ from 0 to $2\pi,$ we obtain the “pure” banana shown in the lower panel of Fig. 12. For this plot, it was assumed that $a=a_{x}^{\prime}=7.3$ kpc (in agreement with the values from Table 1 for $f_{x}=f_{x}^{\prime}=5$ Gyr-1) and $b=c=a/2$. Adopting a nonzero value of $\phi_{0}$ shifts the banana orbit toward an $\infty$-like orbit; for example, the orbit shown in Fig. 5 was obtained with $\phi_{0}=-\pi/5$.
At the end of the first phase of buckling, all orbits can be approximated as banana-like orbits with $f_{z}^{\prime}=2f_{x}$. Using Equation (3), one obtains a relation between the frequencies after buckling:
@@
where $n=(1+\psi)/(2\psi)$ and the values of the parameter $n$ for the selected orbits are given in the last column of Table 1. In order to obtain the new orbits from the Lissajous curves given in the parametric form (5), one needs to replace $m=2,$ which is valid for the banana orbits, with $(2-1/n)$. These values are obviously equal to the $f_{z}^{\prime}/f_{x}^{\prime}$ ratios given in the third column of Table 1. The orbits obtained in this way using the Lissajous curves are shown in Fig. 12. The sizes of the orbits along $x$ were adjusted to the values of $a=a_{x}^{\prime}$ from Table 1, while $b=c=a/2$ and $\phi_{0}=0$ were kept for all of them. It should be noted that in order to close, the orbits are required to cover $4\pi$, $6\pi$, $8\pi$, and $10\pi$ in $\phi$ for $n=2,3,4,$ and 5, respectively.
The origin of these pretzel-like orbits can be understood by recalling that the bending wave speeds up the motion in the horizontal plane. I now focus on the first example of a fish-like orbit shown in the upper panel of Fig. 12. Starting at the upper right corner of the box, at the highest $z$ position and following the orbit, one can see that in comparison with the original banana orbit (bottom panel) the star indeed reaches the next maximum of $z$ not after half a cycle, but much further along the horizontal angle. Overall, the result is that the star makes only three full oscillations along $z$ for each two oscillations in $x$, as expected from the frequency ratio for this orbit, $f_{z}^{\prime}/f_{x}^{\prime}=3/2$. Similar behavior is observed for the next orbits shown in Fig. 12, but to a lesser extent. The orbits with $f_{z}^{\prime}/f_{x}^{\prime}=$ 5/3, 7/4, and 9/5 make five, seven, and nine oscillations in $z$ for each three, four, and five oscillations in $x$, respectively. The limiting case of this sequence is obviously the banana orbit with $f_{z}^{\prime}/f_{x}^{\prime}=2$, which remains unchanged.
The positions of the first four orbits from Table 1 and Fig. 12 in the $a_{x}^{\prime}(a_{x})$ and $f_{x}^{\prime}(f_{x})$ planes are shown as black dots in Fig. 11. The variation of these two quantities in the second phase of buckling suggests that they change adiabatically. As discussed in Section 3.6.2 of Binney & Tremaine (2008), the increase of the horizontal frequency of a harmonic oscillator by a factor $(1+\psi)$ described by Equation (3) should correspond to the decrease of the amplitude of oscillations according to $a_{x}^{\prime}=a_{x}(1+\psi)^{-1/2}$. The numerical values in Table 1 indeed obey these relations.
@@ 6 Discussion
In this work, I revisited the issue of the origin and interpretation of the buckling instability in galactic bars. It was proposed that the phenomenon can be divided into two distinct phases. The first one involves the growth of the distortion of the bar out of the disk plane due to the vertical resonance of stellar orbits in the bar. The growth of the distortion can be described with the mechanism of the driven harmonic oscillator where the initial small perturbation acts as a driving force and modifies the vertical frequencies of other orbits. The distortion retains its orientation with respect to the bar until reaching a maximum value.
@@
As a result of the increased horizontal frequencies, in the second phase of buckling the pattern speed of the vertical distortion is no longer equal to the pattern speed of the bar – it is higher – and the distortion starts to wind up. The banana orbits that dominated the orbital structure of the bar at the time of the resonance evolve into pretzel-like orbits with vertical-to-horizontal frequency ratios below two. The winding up of the distortion also explains why the bar becomes weaker after buckling. Namely, the winding up is associated with faster rotation of the stars in the horizontal plane, which means that the orbits become less radial.
@@
The equation of a driven harmonic oscillator (although without the force growing in time) was previously used in the context of bar buckling by Collier (2020) in an attempt to relate buckling to the phenomenon of Landau damping. Here, instead, this equation was used to demonstrate how the initial small deviation from the vertical symmetry can build up, leading to a strong global distortion of the bar, and the scenario bears no relation to Landau damping.
@@
The violent buckling event seen in the simulation described here, leading to the strong changes in the orbital structure of the bar, seems to be characteristic of pure disks embedded in dark-matter halos, with no bulge component. In a recent study, McClure et al. (2025) looked into the phenomenon using simulations with different bulges of increasing mass. In the case without the bulge, their results are very similar to the ones discussed here; namely, after buckling the orbital frequencies display a very tight relation similar to the one shown in the right panel of Fig. 1. With increasing bulge mass, the resonant orbits still appear, but they do not evolve strongly toward vertical-to-horizontal frequency ratios below two, and they remain close to this resonant value. It seems, therefore, that the bulge does not inhibit the appearance of the vertical resonance, but it does prevent the passage to the second phase of buckling as described here. This is understandable given the importance of mass redistribution in changing the horizontal frequencies, mentioned above. With the additional massive component in the center, this must certainly look different.
@@
Recently, Zozulia et al. (2024) studied the vertical evolution of a mature simulated bar, long after the violent buckling event. They reported behavior similar to that described here, namely the catching of the orbits at resonance and their later passage to vertical-to-horizontal frequency ratios below two, which they refer to as the “circulation mode”. This stage of bar evolution corresponds to the period after $t=5$ Gyr in the simulation discussed here when the distortion is preserved only in the outer part of the bar (last panels of Fig. 13). As discussed in L19, this distortion survives for a few gigayears, shifting toward the outer part of the bar and seeding the second episode of buckling later on. It therefore seems that as the bar grows, its outer parts continuously increase amplitudes of their oscillations in the vertical direction in a kind of quiescent buckling. The process described by Zozulia et al. (2024) is thus not essentially different from the violent buckling event studied here. The latter only happens on a shorter timescale and in the whole inner part of the bar. This picture of the moving distortion also agrees with the early results of Pfenniger & Friedli (1991), which found (their Fig. 14) that the position of the vertical resonance shifts to larger radii after the formation of the BP shape.
@@
A few open questions remain concerning the analysis presented here. For example, it is still not clear why the pattern speed of the vertical distortion increases in such a particular way in the second stage of buckling and becomes equal to a third of the circular frequency. A related question concerns why the banana orbits in the outer part of the bar remain unchanged. This region is characterized by the horizontal frequency equal to twice the pattern speed of the bar, so the pattern speed of the vertical distortion remains equal to the pattern speed of the bar and the distortion does not wind up there.
@@
In summary, this work demonstrates that the phenomenon of buckling instability is not related to the fire-hose instability known from plasma physics, but instead it is triggered by the vertical resonance of bar orbits. The resonance leads to the growing distortion of the bar, which later winds up, resulting in a transformation of banana-like orbits into pretzel-like ones and the formation of a BP shape.
@@ Acknowledgements.
I am grateful to the anonymous referee for useful comments. Computations for this work have been performed using the computer cluster at the Nicolaus Copernicus Astronomical Center of the Polish Academy of Sciences (CAMK PAN).
Appendix A Later evolution of the distortion pattern
This appendix presents the continuation of Fig. 3 and shows the evolution of the vertical distortion pattern of the bar after it starts to wind up.
Figure 13: Same as Fig. 3, but for later times, $t=4.7$ to 5.25 Gyr.